OpenStudy (anonymous):

Solve. [tan(x/Pi-3)-x^3+4x^2+x-4]^4+(4^x-18*2^x+32)^2=0

OpenStudy (asnaseer):

since each term is raised to an even power, then each term must equal zero, so you have two simpler equations to solve:\[\tan(\frac{x}{\pi}-3)-x^3+4x^2+x-4=0\]and:\[4^x-18*2^x+32=0\]

OpenStudy (anonymous):

i drew graph for you

OpenStudy (anonymous):

Sorry, it's tan(x/(Pi-3))

OpenStudy (anonymous):

tan(Pi/(x-3)) . My bad

OpenStudy (asnaseer):

so the two equation to solve are:\[\tan(\frac{\pi}{x-3})−x^3+4x^2+x−4=0\]and:\[4^x-18*2^x+32=0\]

OpenStudy (anonymous):

I think I'll manage with the bottom one, but the first?

OpenStudy (asnaseer):

that's exactly what I was thinking! ;-)

OpenStudy (anonymous):

And how do you solve it?

OpenStudy (anonymous):

tan(Pi/(x-3))-x^3+4x^2+x-4

OpenStudy (anonymous):

spread the love happy eatser y'all

OpenStudy (anonymous):

Suddely a wild troll appears...

OpenStudy (anonymous):

Suddenly*

OpenStudy (asnaseer):

is this for "natural" values of x again?

OpenStudy (anonymous):

Yes, sum is needed in the end.

OpenStudy (asnaseer):

so you should be able to follow the same principals as in the last question you posted.

OpenStudy (anonymous):

All of that must be done again?

OpenStudy (anonymous):

\[\left(\tan(\frac{x}{\pi-3})-x^3+4x^2+x-4\right)^4+(4^x-18\times2^x+32)^2=0\]this is what you are trying to solve????

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

good lord!

OpenStudy (anonymous):

That's what I'm saying.

OpenStudy (anonymous):

second one is easy enough, but how do you do the first?

OpenStudy (anonymous):

And that's what I would love to find out.

OpenStudy (anonymous):

oh i have an idea doh

OpenStudy (anonymous):

we are solving for x right? and of course both have to be equal to zero, since both terms are postive

OpenStudy (anonymous):

but the second term is zero only if x = 1 or x = 4

OpenStudy (asnaseer):

and whatever answers you get MUST make both terms zero, therefore solution to second (simpler) equation will be enough.

OpenStudy (anonymous):

x cannot be one number for the first term and another for the second right? must be the same x either that or i am very confused

OpenStudy (asnaseer):

satellite73: snap!

OpenStudy (anonymous):

and since only 1 and 4 work for second term, either they work for the first term or they do not, so we can check them and see which if any work

OpenStudy (asnaseer):

so just fid solutions of second (simpler) equation and see which one(s) satisfy the first equation as well.

OpenStudy (anonymous):

I like the idea.

OpenStudy (anonymous):

that was what i am thinking

OpenStudy (anonymous):

1 does not seem to work

OpenStudy (asnaseer):

4 is the only working solution.

OpenStudy (anonymous):

let me try that again correctly

OpenStudy (anonymous):

doesn't look like 4 works either

OpenStudy (asnaseer):

satellite73: the "tan" term you have is incorrect - please se correction by ErkoT

OpenStudy (anonymous):

oooooh whew it is \(\tan(\frac{\pi}{x-3})\) that makes it much nicer

OpenStudy (asnaseer):

:)

OpenStudy (anonymous):

but wait, why doesn't one work?

OpenStudy (anonymous):

oh because it is undefined there nvm

OpenStudy (anonymous):

so we check \[\tan(\pi)-4^3+4\times 4^2+4-4=0\] yes? and whew it works

OpenStudy (anonymous):

what a funky problem. a good challenge actually, i will see if i can remember it

OpenStudy (anonymous):

where did it come from?

OpenStudy (anonymous):

Our teacher gave us a test to solve on our own at home. I'm in the eleventh grade at the moment.